# Topology of Discrete Quantum Feedback Control

### v2

@@ V.3 Ten-fold symmetry classification of feedback control with projective measurement
where $\tilde{\mathcal{E}}$ is the matrix representation of $\mathcal{E}$ acting on the doubled Hilbert space (see Sec. III.1), $\tilde{\mathcal{V}}_{X}\ (X=P,C,K,Q)$ is a unitary operator on the doubled Hilbert space, $\epsilon_{X}=\pm 1\ (X=C,K,Q)$, and $\eta_{X}=\pm 1\ (X=C,K)$. Here, we assume that a CPTP map does not have unitary symmetry; if a CPTP map has unitary symmetry, we first block-diagonalize it accoding to the symmetry eigenvalues and consider the BL symmetry class of each block. The combinations of the four types of symmetries (63a)-(63d) lead to 38 classes of non-Hermitian matrices [34].
Whereas general non-Hermitian matrices are classified in terms of the 38 BL symmetry classes [34], some of them are not compatible with CPTP maps for projective measurement. A CPTP map for a projective measurement is given by
@@
where $\hat{P}_{m}$ is a projection operator satisfying $\hat{P}_{m}^{\dagger}=\hat{P}_{m}$ and $\hat{P}_{m}\hat{P}_{m^{\prime}}=\delta_{m,m^{\prime}}\hat{P}_{m}$. A key property of the projective measurement channel (64) is that its eigenvalues are either zero or one, which follows from the Hermiticity and idempotency of projection operators (see Appendix C). Physically, this property is related to the repeatability of projective measurements, that is, repeated projective measurements after the first projective measurement with outcome $m$ always yield the same outcome $m$ [47]. To identify the possible symmetry classes of the projective measurement channel (64), we note that a certain BL symmetry places a constraint on the eigenvalues of a CPTP map. Let $\xi$ be an eigenvalue of a CPTP map $\mathcal{E}$. Then, if this CPTP map has $P$ symmetry (63a) or $C$ symmetry with $\epsilon_{C}=-1$, $-\xi$ is also an eigenvalue of $\mathcal{E}$ [68]. Similarly, if a CPTP map $\mathcal{E}$ has $K$ ($Q$) symmetry, $\epsilon_{K}\xi^{*}$ ($\epsilon_{Q}\xi^{*}$) is also an eigenvalue of $\mathcal{E}$. Since a projective measurement channel (64) cannot have an eigenvalue $-1$, neither $P$ symmetry nor $X=C,K,Q$ symmetry with $\epsilon_{X}=-1$ can be a symmetry of the projective measurement channel (see Fig. 11). Thus, the possible symmetries of a CPTP map for projective measurement are the following ones:
@@
Figure 11: Eigenspectrum of a projective measurement channel (shown by green dots) consisting of eigenvalues zero and one. Since a projective measurement channel cannot have eigenvalue $-1$, its possible BL symmetry classes are restricted to the ten symmetry classes listed in Table 2 which constitute a subset of the entire 38 symmetry classes.
Now let us consider quantum channels for discrete quantum feedback control with projective measurement. Specifically, we consider a CPTP map of the form
@@
The three types of symmetries (65a)-(65c) and the combinations thereof constitute ten symmetry classes of quantum channels for discrete quantum feedback control with projective measurement. Notably, they belong to a subset of the entire 38 BL symmetry classes applicable to general non-Hemitian systems [34, 35]. Each symmetry class is specified by the presence or absence of the three types of symmetries (65a)-(65c) and by the signs of $\eta_{C}$ and $\eta_{K}$, as shown in Table 2. Here, it is sufficient to consider a single symmetry for each type (65a)-(65c). In fact, if a system has multiple symmetries of the same type, we can construct a unitary symmetry by combining the two BL symmetries. We note that the existence of two symmetries among (65a)-(65c) leads to the presence of the remaining symmetry constructed from the combination of the two symmetry operations, resulting in ten classes.
We list the ten-fold symmetry classification of quantum channels in Table 2, where each symmetry class is labeled according to the general classification of non-Hermitian symmetry classes in Ref. [34]. We note that these ten symmetry classes are equivalent to the AZ classes in Ref. [34] if we multiply $i$ to the original CPTP maps. In fact, the $C$ symmetry (65a) and the $K$ symmetry (65b) are equivalent to “time-reversal symmetry” and “particle-hole symmetry” of $i\tilde{\mathcal{E}}$, respectively [97, 34]. In Table 2, we also show the AZ classes corresponding to the symmetry classes of CPTP maps.
@@
Even if a measurement is not projective, other symmetry classes are still not allowed unless eigenvalues of a CPTP map for measurement always appear as pairs like $(\xi,-\xi)$ or $(\xi,-\xi^{*})$ [see the paragraph above Eqs. (65a)-(65c)]. For example, if a measurement is not projective due to an error in the measurement process, the possible symmetry classes are not altered because we can take an error-free limit under which the symmetry is intact.
The ten AZ symmetry classes are formally equivalent to the symmetry classes for generators of Markovian quantum dynamics discussed in Ref. [65]. In fact, the symmetry classes that are given in Table 2 agree with the symmetry classes of CPTP maps that are generated by Lindbladians (see Appendix D). However, we note that a CPTP map for feedback control cannot have a generator if it has the zero eigenvalue [see Eq. (39)]. In general, a CPTP map generated by a Lindbladian can be classified by the 38 BL symmetry classes if one includes shifted symmetries [98, 66, 67]. However, a projective measurement channel (64) cannot have those shifted symmetries because the number of zero eigenvalues is different from that of unit eigenvalues, while shifted symmetry is accompanied by the symmetry of the eigenspectrum with respect to a certain line [98, 66, 67]. Thus, we conclude that the ten-fold symmetry classification exhausts all relevant symmetry classes of discrete quantum feedback control with projective measurement.
We note that a CPTP map always has the modular conjugation symmetry [see Eq. (62) in Sec. V.2]. The modular conjugation symmetry corresponds to a $K$ symmetry with $\eta_{K}=+1$. Thus, a quantum channel as a whole always has this symmetry. However, if a quantum channel possesses an additional unitary symmetry, special care must be taken. This is because the quantum channel can then be block-diagonalized according to the symmetry eigenvalues, and each symmetry sector (i.e. each block) does not necessarily have to have the modular conjugation symmetry. In this case, we must consider an appropriate BL symmetry class in each sector. We will revisit this point when we discuss concrete examples in Sec. VI.
Here, we provide sufficient conditions for the BL symmetries of CPTP maps to satisfy Eqs. (65a)-(65c), which are helpful to gain an insight into feedback control with symmetry. The symmetries in Eqs. (65a), (65b), and (65c) are satisfied if Kraus operators satisfy
@@
respectively, where $\hat{K}_{m}\ (m=1,2,\cdots,N_{\mathrm{K}})$ are the Kraus operators, $N_{\mathrm{K}}$ is the number of the Kraus operators, and $f$ is a bijection on $\{1,2,\cdots,N_{\mathrm{K}}\}$. Suppose that the Kraus operators of feedback control are given by $\hat{K}_{m}=\hat{U}_{m}\hat{M}_{m}$, where $\hat{M}_{m}$ is a measurement operator and $\hat{U}_{m}=e^{-i\hat{H}_{m}\tau}$ is a unitary operator conditioned on a measurement outcome $m$. Then, if the two conditions
@@
In contrast, the symmetries (65a) and (65c) respectively include transposition and Hermitian conjugation, both of which reverse the order of operators. Because of this order-reversing nature, the symmetries (65a) and (65c) are unlikely to be realized with the simplest protocol involving a single measurement and feedback since $(\hat{U}_{m}\hat{M}_{m})^{T}=\hat{M}_{m}^{T}\hat{U}_{m}^{T}$ and $(\hat{U}_{m}\hat{M}_{m})^{\dagger}=\hat{M}_{m}^{\dagger}\hat{U}_{m}^{\dagger}$. We will demonstrate in Sec. VI that those symmetries can be satisfied by a protocol with a two-point measurement before and after feedback control, where the Kraus operators are given by $\hat{K}_{m}=\hat{M}_{m_{2}}\hat{U}_{m_{1}}\hat{M}_{m_{1}}$ with $m=(m_{1},m_{2})$ being a pair of measurement outcomes (see Sec. II).
@@ VI Feedback control with symmetry
In this section, we present examples of feedback control for each symmetry class shown in Table 2. On the basis of the intrinsic modular conjugation symmetry (see Sec. V.2), the ten symmetry classes of quantum channels listed in Table 2 can be categorized into three groups according the presence or absense of $K$ symmetry and the sign of $\eta_{K}$. The first group (AI, AI $+$ psH+, and AI $+$ psH-) includes $K$ symmetry with $\eta_{K}=+1$. Since the modular conjugation symmetry features $K$ symmetry with $\eta_{K}=+1$, these symmetry classes can be realized with quantum channels without additional symmetries or with additional $C$ or $Q$ symmetry. The second group (A, psH, AI, and AII) does not have $K$ symmetry. These classes can be realized in a symmetry sector of a quantun channel with unitary symmetry. The third group (AII, AII $+$ psH+, and AII $+$ psH-) includes $K$ symmetry with $\eta_{K}=-1$. Such symmetry classes can be realized by imposing $K$ symmetry with $\eta_{K}=-1$ on the second group.
@@ VI.1 Class AI: feedback without symmetry or with particle-hole symmetry
We begin with class AI, which includes only $K$ symmetry with $\eta_{K}=+1$. Since a CPTP map always has the modular conjugation symmetry (see Sec. V.2), a quantum channel belongs to class AI if it does not have any other symmetry. For example, the chiral Maxwell’s demon presented in Sec. IV belongs to this class.
A CPTP map can have $K$ symmetry with $\eta_{K}=+1$ that is not equivalent to the modular conjugation symmetry. For example, if the measurement operators and the unitary operators for feedback control satisfy Eq. (69), the corresponding CPTP map has $K$ symmetry. Since the first condition in Eq. (69) follows from the particle-hole symmetry of the feedback Hamiltonian, this symmetry is physically interperted as the particle-hole symmetry in feedback control. If a CPTP map has $K$ symmetry (65b) that is not equivalent to the modular conjugation symmetry, we can construct a unitary symmetry
@@
The CPTP map can be block-diagonalized by using the unitary symmetry (71), and the $K$ symmetry and the modular conjugation symmetry are equivalent in each symmetry sector. Thus, if the CPTP map does not have any other symmetry, the block-diagonalized CPTP map belongs to class AI.
VI.2 Class AI $+$ psH+ and AI $+$ psH-: feedback with a two-point projective measurement
The symmetry classes that include $K$ symmetry with $\eta_{K}=+1$ are class AI, AI $+$ psH+, and AI $+$ psH-, where the last two classes include $C$ and $Q$ symmetries. As mentioned in Sec. V, $C$ and $Q$ symmetries respectively involve transposition and Hermitian conjugation, both of which reverse the order of operators. Here we construct quantum channels with $C$ and $Q$ symmetries by using feedback control with two-point projective measurements. Specifically, we consider a scheme with two projective measurements intervened by a unitary operation. This process is described by a CPTP map
@@ VI.2 Class AI $+$ psH+ and AI $+$ psH-: feedback with a two-point projective measurement
where $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1}):=|(\hat{U}_{\vec{j}_{1},a_{1}})_{\vec{j}_{2},a_{2};\vec{j}_{1},a_{1}}|^{2}$ is the probability of a measurement outcome $(\vec{j}_{2},a_{2})$ being obtained in the projective measurement performed after feedback control associated with a measurment outcome $(\vec{j}_{1},a_{1})$. Thus, the conditions (68a) and (68c) for $C$ and $Q$ symmetries follow from symmetry of the transition probability $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1})$.
@@
is the truncated Bloch matrix. Nonzero eigenvalues of the Bloch matrix are completely determined by those of the truncated Bloch matrix. The BL symmetry class of the Bloch matrix is also determined from that of the truncated Bloch matrix.
@@
Let us examine the symmetry of the truncated Bloch matrix. First, we note that the truncated Bloch matrix always has an antiunitary symmetry (the modular conjugation symmetry)
$X_{\mathrm{trunc}}^{*}(-\vec{k})=X_{\mathrm{trunc}}(\vec{k})$ (82)
due to the Hermiticity-preserving nature of the CPTP map (see Sec. V.2 and Appendix A.4). This corresponds to $K$ symmetry (65b) with $\eta_{K}=+1$. Next, we consider $C$ symmetry (65a). The simplest symmetry of this type with $\eta_{C}=+1$ is given by
@@
which can be interpreted as reciprocity in the feedback process. The truncated Bloch matrix (and hence the original Bloch matrix) with Eq. (83) belongs to the symmetry class AI $+$ psH+ in Table 2. In the present case with Eqs. (82) and (83), the truncated Bloch matrix is Hermitian ($X_{\mathrm{trunc}}^{\dagger}(\vec{k})=X_{\mathrm{trunc}}(\vec{k})$), and therefore its eigenvalues are real. More specifically, the truncated Bloch matrix in a 1D system cannot have a nonzero winding number.
@@
which represents the reciprocity in which the the two spin components flow in opposite directions. In contrast, since $p(\vec{j}_{2},\sigma|\vec{j}_{1},\sigma^{\prime})\geq 0$, conditions (87b) and (87c) are satisfied if and only if
@@
This means that the truncated Bloch matrix is written as
@@
This matrix actually possesses a higher symmetry; the two spin sectors are completely decoupled and the probability in each sector is separately conserved. This is because the Bloch matrix cannot be an arbitrary non-Hermitian matrix; it must respect the complete positivity and the trace-preserving property of the original CPTP map. This implies that the CPTP properties impose nontrivial constraints on the symmetry of CPTP maps, which may also have an impact on the topological classification.
VI.3 Class A, AI, AII, and psH: feedback with unitary symmetry

### v3

@@ V.3 Tenfold symmetry classification of feedback control with projective measurement
where $\tilde{\mathcal{E}}$ is the matrix representation of $\mathcal{E}$ acting on the doubled Hilbert space (see Sec. III.1), $\tilde{\mathcal{V}}_{X}\ (X=P,C,K,Q)$ is a unitary operator on the doubled Hilbert space, $\epsilon_{X}=\pm 1\ (X=C,K,Q)$, and $\eta_{X}=\pm 1\ (X=C,K)$. Here, we assume that a CPTP map does not have unitary symmetry; if a CPTP map has unitary symmetry, we first block diagonalize it according to the symmetry eigenvalues and consider the BL symmetry class of each block. The combinations of the four types of symmetries, Eqs. (63a)-(63d), lead to 38 classes of non-Hermitian matrices [34].
Whereas general non-Hermitian matrices are classified in terms of the 38 BL symmetry classes [34], some of them are not compatible with CPTP maps for projective measurements. A CPTP map for a projective measurement is given by
@@
where $\hat{P}_{m}$ is a projection operator satisfying $\hat{P}_{m}^{\dagger}=\hat{P}_{m}$ and $\hat{P}_{m}\hat{P}_{m^{\prime}}=\delta_{m,m^{\prime}}\hat{P}_{m}$. A key property of the projective measurement channel (64) is that its eigenvalues are either zero or one, which follows from the Hermiticity and idempotency of projection operators (see Appendix C). Physically, this property is related to the repeatability of projective measurements; that is, repeated projective measurements after the first projective measurement with outcome $m$ always yield the same outcome $m$ [47]. To identify the possible symmetry classes of the projective measurement channel (64), we note that a certain BL symmetry places a constraint on the eigenvalues of a CPTP map. Let $\xi$ be an eigenvalue of a CPTP map $\mathcal{E}$. Then, if this CPTP map has $P$ symmetry (63a) or $C$ symmetry with $\epsilon_{C}=-1$, $-\xi$ is also an eigenvalue of $\mathcal{E}$ [68]. Similarly, if a CPTP map $\mathcal{E}$ has $K$ ($Q$) symmetry, $\epsilon_{K}\xi^{*}$ ($\epsilon_{Q}\xi^{*}$) is also an eigenvalue of $\mathcal{E}$. Since a projective measurement channel (64) cannot have an eigenvalue $-1$, neither $P$ symmetry nor $X=C,K,Q$ symmetry with $\epsilon_{X}=-1$ can be a symmetry of the projective measurement channel (see Fig. 11). Thus, the possible symmetries of a CPTP map for the projective measurement are as follows:
@@
Figure 11: Eigenspectrum of a projective measurement channel (shown by green dots) consisting of eigenvalues zero and one. Since a projective measurement channel cannot have eigenvalue $-1$, its possible BL symmetry classes are restricted to the ten symmetry classes listed in Table 2, which constitute a subset of the entire 38 symmetry classes.
Now, let us consider quantum channels for discrete quantum feedback control with projective measurements. Specifically, we consider a CPTP map of the form
@@
The three types of symmetries, Eqs. (65a)-(65c), and the combinations thereof constitute ten symmetry classes of quantum channels for discrete quantum feedback control with projective measurements. Notably, they belong to a subset of the entire 38 BL symmetry classes applicable to general non-Hermitian systems [34, 35]. Each symmetry class is specified by the presence or absence of the three types of symmetries, Eqs. (65a)-(65c), and by the signs of $\eta_{C}$ and $\eta_{K}$, as shown in Table 2. Here, it is sufficient to consider a single symmetry for each type [Eqs. (65a)-(65c)]. In fact, if a system has multiple symmetries of the same type, we can construct a unitary symmetry by combining the two BL symmetries. We note that the existence of two symmetries among Eqs. (65a)-(65c) leads to the presence of the remaining symmetry constructed from the combination of the two symmetry operations, resulting in ten classes.
We list the tenfold symmetry classification of quantum channels in Table 2, where each symmetry class is labeled according to the general classification of non-Hermitian symmetry classes in Ref. [34]. We note that these ten symmetry classes are equivalent to the AZ classes in Ref. [34] if we multiply the original CPTP maps by $i$. In fact, the $C$ symmetry (65a) and the $K$ symmetry (65b) are equivalent to “time-reversal symmetry” and “particle-hole symmetry” of $i\tilde{\mathcal{E}}$, respectively [97, 34]. In Table 2, we also show the AZ classes corresponding to the symmetry classes of CPTP maps.
@@
Even if a measurement is not projective, other symmetry classes are still not allowed unless eigenvalues of a CPTP map for measurements always appear as pairs like $(\xi,-\xi)$ or $(\xi,-\xi^{*})$ [see the paragraph above Eqs. (65a)-(65c)]. For example, if a measurement is not projective due to an error in the measurement process, the possible symmetry classes are not altered because we can take an error-free limit under which the symmetry is intact.
The ten AZ symmetry classes are formally equivalent to the symmetry classes for generators of Markovian quantum dynamics discussed in Ref. [65]. In fact, the symmetry classes that are given in Table 2 agree with the symmetry classes of CPTP maps that are generated by Lindbladians (see Appendix D). However, we note that a CPTP map for feedback control cannot have a generator if it has a zero eigenvalue [see Eq. (39)]. In general, a CPTP map generated by a Lindbladian can be classified by the 38 BL symmetry classes if one includes shifted symmetries [98, 66, 67]. However, a projective measurement channel (64) cannot have those shifted symmetries because the number of zero eigenvalues is different from that of unit eigenvalues, while shifted symmetry is accompanied by the symmetry of the eigenspectrum with respect to a certain line [98, 66, 67]. Thus, we conclude that the tenfold symmetry classification exhausts all relevant symmetry classes of discrete quantum feedback control with projective measurements.
We note that a CPTP map always has modular conjugation symmetry [see Eq. (62) in Sec. V.2]. The modular conjugation symmetry corresponds to a $K$ symmetry with $\eta_{K}=+1$. Thus, a quantum channel as a whole always has this symmetry. However, if a quantum channel possesses an additional unitary symmetry, special care must be taken because the quantum channel can then be block diagonalized according to the symmetry eigenvalues; each symmetry sector (i.e. each block) does not necessarily have modular conjugation symmetry. In this case, we must consider an appropriate BL symmetry class in each sector. We will revisit this point when we discuss concrete examples in Sec. VI.
Here, we provide sufficient conditions for the BL symmetries of CPTP maps to satisfy Eqs. (65a)-(65c), which are helpful to gain insight into feedback control with symmetry. The symmetries in Eqs. (65a)-(65c) are satisfied if Kraus operators satisfy
@@
respectively, where $\hat{K}_{m}\ (m=1,2,\cdots,N_{\mathrm{K}})$ are the Kraus operators, $N_{\mathrm{K}}$ is the number of Kraus operators, and $f$ is a bijection on $\{1,2,\cdots,N_{\mathrm{K}}\}$. Suppose that the Kraus operators of feedback control are given by $\hat{K}_{m}=\hat{U}_{m}\hat{M}_{m}$, where $\hat{M}_{m}$ is a measurement operator and $\hat{U}_{m}=e^{-i\hat{H}_{m}\tau}$ is a unitary operator conditioned on a measurement outcome $m$. Then, if the two conditions
@@
In contrast, the symmetries (65a) and (65c), respectively, include transposition and Hermitian conjugation, both of which reverse the order of operators. Because of this order-reversing nature, the symmetries (65a) and (65c) are unlikely to be realized with the simplest protocol involving a single measurement and feedback since $(\hat{U}_{m}\hat{M}_{m})^{T}=\hat{M}_{m}^{T}\hat{U}_{m}^{T}$ and $(\hat{U}_{m}\hat{M}_{m})^{\dagger}=\hat{M}_{m}^{\dagger}\hat{U}_{m}^{\dagger}$. In Sec. VI, we demonstrate that those symmetries can be satisfied by a protocol with a two-point measurement before and after feedback control, where the Kraus operators are given by $\hat{K}_{m}=\hat{M}_{m_{2}}\hat{U}_{m_{1}}\hat{M}_{m_{1}}$ with $m=(m_{1},m_{2})$ being a pair of measurement outcomes (see Sec. II).
@@ VI Feedback control with symmetry
In this section, we present examples of feedback control for each symmetry class shown in Table 2. On the basis of the intrinsic modular conjugation symmetry (see Sec. V.2), the ten symmetry classes of quantum channels listed in Table 2 can be categorized into three groups according the presence or absence of $K$ symmetry and the sign of $\eta_{K}$. The first group (AI, AI $+$ psH+, and AI $+$ psH-) includes $K$ symmetry with $\eta_{K}=+1$. Since the modular conjugation symmetry features $K$ symmetry with $\eta_{K}=+1$, these symmetry classes can be realized with quantum channels without additional symmetries or with additional $C$ or $Q$ symmetry. The second group (A, psH, AI, and AII) does not have $K$ symmetry. These classes can be realized in a symmetry sector of a quantum channel with unitary symmetry. The third group (AII, AII $+$ psH+, and AII $+$ psH-) includes $K$ symmetry with $\eta_{K}=-1$. Such symmetry classes can be realized by imposing $K$ symmetry with $\eta_{K}=-1$ on the second group.
@@ VI.1 Class AI: Feedback without symmetry or with particle-hole symmetry
We begin with class AI, which includes only $K$ symmetry with $\eta_{K}=+1$. Since a CPTP map always has modular conjugation symmetry (see Sec. V.2), a quantum channel belongs to class AI if it does not have any other symmetry. For example, the chiral Maxwell demon presented in Sec. IV belongs to this class.
A CPTP map can have $K$ symmetry with $\eta_{K}=+1$ that is not equivalent to the modular conjugation symmetry. For example, if the measurement operators and the unitary operators for feedback control satisfy Eq. (69), the corresponding CPTP map has $K$ symmetry. Since the first condition in Eq. (69) follows from the particle-hole symmetry of the feedback Hamiltonian, this symmetry is physically interpreted as the particle-hole symmetry in feedback control. If a CPTP map has $K$ symmetry (65b) that is not equivalent to the modular conjugation symmetry, we can construct a unitary symmetry
@@
The CPTP map can be block diagonalized by using the unitary symmetry (71), and the $K$ symmetry and modular conjugation symmetry are equivalent in each symmetry sector. Thus, if the CPTP map does not have any other symmetry, the block diagonalized CPTP map belongs to class AI.
VI.2 Classes AI $+$ psH+ and AI $+$ psH-: Feedback with a two-point projective measurement
The symmetry classes that include $K$ symmetry with $\eta_{K}=+1$ are classes AI, AI $+$ psH+, and AI $+$ psH-, where the last two classes include $C$ and $Q$ symmetries. As mentioned in Sec. V, $C$ and $Q$ symmetries, respectively, involve transposition and Hermitian conjugation, both of which reverse the order of operators. Here we construct quantum channels with $C$ and $Q$ symmetries by using feedback control with two-point projective measurements. Specifically, we consider a scheme with two projective measurements intervened by a unitary operation. This process is described by a CPTP map
@@ VI.2 Classes AI $+$ psH+ and AI $+$ psH-: Feedback with a two-point projective measurement
where $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1}):=|(\hat{U}_{\vec{j}_{1},a_{1}})_{\vec{j}_{2},a_{2};\vec{j}_{1},a_{1}}|^{2}$ is the probability of a measurement outcome $(\vec{j}_{2},a_{2})$ being obtained in the projective measurement performed after feedback control associated with a measurement outcome $(\vec{j}_{1},a_{1})$. Thus, the conditions (68a) and (68c) for $C$ and $Q$ symmetries follow from symmetry of the transition probability $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1})$.
@@
is the truncated Bloch matrix. Nonzero eigenvalues of the Bloch matrix are completely determined by those of the truncated Bloch matrix. The BL symmetry class of a given Bloch matrix is also determined from that of the corresponding truncated Bloch matrix.
@@
Let us examine the symmetry of the truncated Bloch matrix. First, we note that the truncated Bloch matrix always has an antiunitary symmetry (the modular conjugation symmetry),
$X_{\mathrm{trunc}}^{*}(-\vec{k})=X_{\mathrm{trunc}}(\vec{k}),$ (82)
due to the Hermiticity-preserving nature of the CPTP map (see Sec. V.2 and Appendix A.4). This case corresponds to $K$ symmetry (65b) with $\eta_{K}=+1$. Next, we consider $C$ symmetry (65a). The simplest symmetry of this type with $\eta_{C}=+1$ is given by
@@
which can be interpreted as reciprocity in the feedback process. The truncated Bloch matrix (and hence the original Bloch matrix) with Eq. (83) belongs to the symmetry class AI $+$ psH+ in Table 2. In the present case with Eqs. (82) and (83), the truncated Bloch matrix is Hermitian [$X_{\mathrm{trunc}}^{\dagger}(\vec{k})=X_{\mathrm{trunc}}(\vec{k})$], and therefore its eigenvalues are real. More specifically, the truncated Bloch matrix in a 1D system cannot have a nonzero winding number.
@@
which represents the reciprocity in which the two spin components flow in opposite directions. In contrast, since $p(\vec{j}_{2},\sigma|\vec{j}_{1},\sigma^{\prime})\geq 0$, conditions (87b) and (87c) are satisfied if and only if
@@
Thus, the truncated Bloch matrix is written as
@@
This matrix actually possesses a higher symmetry; the two spin sectors are completely decoupled, and the probability in each sector is separately conserved. The Bloch matrix cannot be an arbitrary non-Hermitian matrix; it must respect the complete positivity and the trace-preserving property of the original CPTP map. Thus, the CPTP properties impose nontrivial constraints on the symmetry of CPTP maps, which may also have an impact on the topological classification.
VI.3 Classes A, AI, AII, and psH: Feedback with unitary symmetry